# WebRL: Training LLM Web Agents via Self-Evolving Online Curriculum Reinforcement Learning

### v1

@@ 3.9 Case Study
Figure 8 presents some instructions generated by the self-evolving curriculum learning strategy across different phases. Although these instructions are grouped by phase, the instructions shown in phase $i+1$ are not necessarily generated using the instructions from phase $i$ as seeds. As the phase increases, two types of augmentation occur for previously incomplete instructions. In one case, new instructions with similar task requirements are generated, enabling the model to complete the previously unfinishable instructions after working through these newly generated ones. For example, In phase 2, the instruction improves upon the phase 1 instruction by offering a clearer task description and explicitly requiring results in “yearly interval”. This enhancement enables the model to complete the task successfully by removing ambiguity. Additionally, with the positive feedback from the clarified phase 2 instruction, the model can better understand and accurately perform the original Phase 1 task as well. In another case, the model generates tasks with increased complexity and greater diversity, which the agent is unable to complete successfully. This task complication facilitates a continuous improvement in the model’s capabilities by challenging its performance boundaries. Therefore, the process of instruction generation exhibits such a pattern: for tasks that the model is unable to perform, analogous tasks are created to provide incremental steps that facilitate learning how to accomplish that type of task. Furthermore, tasks that remain challenging for the model are also generated and undergo the aforementioned iterative process. Through this iterative approach, the model’s capabilities gradually improve, enabling it to perform increasingly complex tasks over time.
@@ 4 Related Works
Adopting LLMs as Agent. As LLMs capabilities advance, their applications extend beyond text generation (Zheng et al., 2023) and complex reasoning (Zelikman et al., 2024; Zhang et al., 2024d), and increasingly involve acting as agents for device control. Current research in this area falls into two main categories: training-free and training-based approaches. Training-free methods enhance pre-existing LLMs through prompt engineering (Yan et al., 2023; He et al., 2024; Zhang et al., 2024c) and constructing complex systems (Liu et al., 2023b; Wang et al., 2023; Yang et al., 2023; Wu et al., 2024; Iong et al., 2024; Zhang et al., 2024a). However, their performance is constrained by the limitations of the underlying LLMs, and the absence of fine-tuning restricts further improvement (Chen et al., 2023; Zeng et al., 2023; Xie et al., 2024). Training-based approaches, primarily relying on imitation learning, require extensive expert demonstrations (Gur et al., 2023; Zhang & Zhang, 2023; Deng et al., 2024; Hong et al., 2024; Rawles et al., 2024; Zhang et al., 2024b), which are costly to obtain. Although some methods use powerful LLMs like GPT-4 to generate demonstrations (Chen et al., 2023), their accuracy remains insufficient for complex tasks. These methods often maximize the likelihood of individual actions without adequately considering long-term effects, limiting generalization (Ghosh et al., 2021; Bai et al., 2024). To mitigate this, some studies use sampling-based methods to estimate long-term effects (Lai et al., 2024; Putta et al., 2024), while others, like ours, leverage reinforcement learning (Bai et al., 2024; Pan et al., 2024; Zhai et al., 2024). However, most existing methods rely on static task sets, which hinder the agent’s continuous improvement as its capabilities evolve. To overcome this, we propose a dynamic task generation framework that adjusts task complexity based on the agent’s progress, alongside a KL-constrained policy-update algorithm for ongoing performance enhancement.
@@
Reinforcement Learning for LLMs. Reinforcement learning has gained traction in training LLMs, with applications ranging from preference optimization (Ouyang et al., 2022; Casper et al., 2023) to complex reasoning (Hao et al., 2023; Pang et al., 2024). A growing area of interest involves using RL for device control tasks, which require multi-step interactions where the model selects appropriate actions based on the device state. This sequential decision-making aligns well with RL techniques. Existing research has explored RL-trained LLM agents for complex device control, primarily using online learning methods. For instance, AgentQ (Putta et al., 2024) uses DPO (Rafailov et al., 2024b) for policy updates based on interaction data, while other methods (Bai et al., 2024; Zhou et al., 2024; Zhai et al., 2024) utilize actor-critic architectures, which we also adopt. However, in web tasks, feedback is often limited to binary success or failure after multiple interaction rounds. This can penalize correct intermediate actions due to later mistakes, complicating the reuse of previous data. Moreover, current research mainly applies reinforcement learning techniques on predefined fixed task sets, far from fully exploring its potential for continuous improvement through trial and error. To address this, we propose an autonomous curriculum learning mechanism that dynamically generates tasks based on the agent’s evolving skills, fostering ongoing progress. Additionally, we introduce a KL-constrained policy update algorithm and a specialized replay buffer to reuse valuable historical data and prevent knowledge forgetting during iterative curriculum updates.
@@ 5 Conclusion
Appendix A Training Details
WebRL and baselines: For RL-based baselines (except DigiRL), the interaction data from WebRL’s first phase is used, while DigiRL is trained using the first-phase instructions in an online learning setup. Hence, except for DigiRL, the other RL baselines fall under offline reinforcement learning. We reproduce the same framework used in DigiRL within the WebArena environment. Specifically, we use the same components, including the (AWR) method, the instruction-level and step-level value functions, and the replay buffer described in DigiRL. The main modifications we make are adjusting the data format to align with /model and tweaking certain hyperparameters.
In each phase of learning within WebRL, 500 new instructions that meet the filtering criteria are selected from those generated by GPT-4o. Both newly generated interaction data and historical data with perplexity between 1/0.95 and 1/0.5 from the replay buffer are used to train the actor and critic, with the replay data limited to twice the size of current interaction data. The hyperparameters employed in WebRL and all baselines are presented in Table 4.

### v2

@@ 3.9 Case Study
Figure 8 presents some instructions generated by the self-evolving curriculum learning strategy across different phases. Although these instructions are grouped by phase, the instructions shown in phase $i+1$ are not necessarily generated using the instructions from phase $i$ as seeds. As the phase increases, two types of augmentation occur for previously incomplete instructions. In one case, new instructions with similar task requirements or lower difficulty levels are generated. These new instructions help the model to successfully complete previously unfinishable tasks by allowing it to practice and build competence with the easier or related tasks first. For example, in phase 2, the instruction improves upon the phase 1 instruction by offering a clearer task description and explicitly requiring results in “yearly interval”. This enhancement enables the model to complete the task successfully by removing ambiguity. Additionally, changing the description of the instruction can improve agent exploration. With the positive feedback from the clarified phase 2 instruction, the model can better understand and accurately perform the original phase 1 task as well. In another case, tasks with increased complexity and diversity are generated. This task complication facilitates a continuous improvement in the model’s capabilities by challenging its performance boundaries. Therefore, the process of instruction generation exhibits such a pattern: for tasks that the model is unable to perform, analogous tasks are created to provide incremental steps that facilitate learning how to accomplish that type of task. Furthermore, tasks that remain challenging for the model are also generated and undergo the aforementioned iterative process. Through this cycle of challenge and refinement, the model’s capabilities gradually expand, enabling it to handle increasingly complex tasks over time.
@@ 4 Related Works
Adopting LLMs as Agent. As LLM capabilities advance, their applications extend beyond text generation (Zheng et al., 2023; Zhao et al., 2023) and complex reasoning (Zelikman et al., 2024; Zhang et al., 2024d), and increasingly involve acting as agents for device control. Current research in this area falls into two main categories: training-free and training-based approaches. Training-free methods enhance pre-existing LLMs through prompt engineering (Yan et al., 2023; He et al., 2024; Zhang et al., 2024c) and constructing complex systems (Liu et al., 2023; Yang et al., 2023; Wang et al., 2024a; Wu et al., 2024; Iong et al., 2024; Zhang et al., 2024a). However, their performance is constrained by the limitations of the underlying LLMs, and the absence of fine-tuning restricts further improvement (Chen et al., 2023; Zeng et al., 2023; Xie et al., 2024). Training-based approaches, primarily relying on imitation learning, require extensive expert demonstrations (Zhang & Zhang, 2023; Gur et al., 2024; Deng et al., 2024; Hong et al., 2024; Rawles et al., 2024; Zhang et al., 2024b), which are costly to obtain. Although some methods use powerful LLMs like GPT-4 to generate demonstrations (Chen et al., 2023), their accuracy remains insufficient for complex tasks. These methods often maximize the likelihood of individual actions without adequately considering long-term effects, limiting generalization (Ghosh et al., 2021; Bai et al., 2024). To mitigate this, some studies use sampling-based methods to estimate long-term effects (Lai et al., 2024; Putta et al., 2024), while others, like ours, leverage reinforcement learning (Carta et al., 2023; Bai et al., 2024; Pan et al., 2024; Tan et al., 2024; Zhai et al., 2024). However, most existing methods rely on static task sets, which hinder the agent’s continuous improvement as its capabilities evolve. To overcome this, we propose a dynamic task generation framework that adjusts task complexity based on the agent’s progress, alongside a policy-update algorithm for ongoing performance enhancement.
@@
Reinforcement Learning for LLMs. Reinforcement learning (RL) has gained traction in training large language models (LLMs), with applications ranging from preference optimization (Ouyang et al., 2022; Casper et al., 2024) to complex reasoning (Hao et al., 2023; Pang et al., 2024). A growing area of interest involves using RL for device control tasks, which require multi-step interactions where the model selects appropriate actions based on the device state. This sequential decision-making aligns well with RL techniques. Existing research has explored RL-trained LLM agents for complex device control, primarily using online learning methods. For instance, AgentQ (Putta et al., 2024) uses DPO (Rafailov et al., 2024b) for policy updates based on interaction data, while other methods (Bai et al., 2024; Zhou et al., 2024b; Zhai et al., 2024) utilize actor-critic architectures, which we also adopt. However, in web tasks, feedback is often limited to binary success or failure after multiple interaction rounds. This can penalize correct intermediate actions due to later mistakes, complicating the reuse of previous data. Moreover, current research tends to focus on a fixed set of tasks for comparison with imitation learning, limiting the potential for continuous improvement through trial and error. To address this, we propose an autonomous curriculum learning mechanism that dynamically generates tasks based on the agent’s evolving skills, fostering ongoing progress. Additionally, we introduce a KL-constrained policy update algorithm and a specialized replay buffer to reuse valuable historical data and prevent knowledge forgetting during iterative curriculum updates.
@@ 5 Conclusion
Appendix A Details of Policy Update Algorithm in WebRL
A.1 Derivation
By substituting eq. 3 into eq. 2, we can obtain:
$\pi^{*}(a_{t}|s_{t},I)=\pi_{\text{ref}}(a_{t}|s_{t},I)\exp\left(\frac{1}{\beta}A^{*}(s_{t},a_{t},I)\right)$ (10)
$\pi^{*}$ is the optimal solution to our target (eq. 1). If $\pi_{\theta}$ is represented by a function approximator, we aim to make it as close as possible to $\pi^{*}$.
$\begin{split}&\operatorname*{arg\,min}\limits_{\theta}\mathbb{E}_{\nu}\left[\left(\log\pi_{\theta}(a|s,I)-\log\pi^{*}(a|s,I)\right)^{2}\right]\\ =&\operatorname*{arg\,min}\limits_{\theta}\mathbb{E}_{\nu}\left[\left(\log\pi_{\theta}(a|s,I)-\log\pi_{\text{ref}}(a|s,I)-\frac{1}{\beta}A^{*}(s,a,I)\right)^{2}\right]\\ =&\operatorname*{arg\,min}\limits_{\theta}\mathbb{E}_{\nu}\left[\left(\beta\log\frac{\pi_{\theta}(a|s,I)}{\pi_{\text{ref}}(a|s,I)}-A^{*}(s,a,I)\right)^{2}\right]\end{split}$ (11)
Hence, the loss function can be defined as:
where $\nu$ represents the distribution of experience used for training. For any given state-action pair, $\pi_{\theta}$ is expected to match the target policy $\pi^{*}(a|s,I)$. There is no restriction on the data distribution $\nu$, indicating the algorithm can function effectively in an off-policy setting. Since $A*$ is unknown, we estimate $A^{*}(s,a,I)$ using eq. 7 and eq. 8 to train the policy $\pi_{\theta}$. The feasibility of this approach is demonstrated in Appendix A.2.
Further Analysis. Although $\nu$ can follow any distribution, achieving stable policy improvement often requires some control over $\nu$. The primary goal is to enhance the probability of actions that successfully complete the task and to address the deficiencies of the current policy $\pi_{\theta}$. Therefore, $\nu$ typically consists of data sampled from the current policy being trained and prior successful experiences.
Why not use KL divergence to measure the distance between $\pi^{*}$ and $\pi_{\theta}$. For eq. 11, many studies use KL divergence to measure the distance between two policies. When KL divergence is used, the optimization goal of $\theta$ is:
$\begin{split}&\operatorname*{arg\,min}\limits_{\theta}\mathbb{E}_{s\sim d(s)}\left[D_{\text{KL}}\left(\pi^{*}(\cdot|s,I)||\pi_{\theta}(\cdot|s,I)\right)\right]\\ =&\operatorname*{arg\,max}\limits_{\theta}\mathbb{E}_{s\sim d(s)}\mathbb{E}_{a\sim\pi^{*}(a|s,I)}\left[\log\pi_{\theta}(a|s,I)\right]\\ =&\operatorname*{arg\,max}\limits_{\theta}\mathbb{E}_{s\sim d(s)}\int_{a}\pi_{\text{ref}}(a|s,I)\exp(\frac{1}{\beta}A^{*}(s,a,I)\log\pi_{\theta}(a|s,I)\text{d}a\\ =&\operatorname*{arg\,max}\limits_{\theta}\mathbb{E}_{s\sim d(s)}\mathbb{E}_{a\sim\pi_{\text{ref}}(a|s,I)}\left[\log\pi_{\theta}(a|s,I)\exp(\frac{1}{\beta}A^{*}(s,a,I))\right]\end{split}$ (13)
where $d(s)$ is a distribution of state. The choice of using eq. 11 instead of eq. 13 as the optimization target can be explained as follows: (1) Eq. 13 imposes stronger restrictions on the training data. Specifically, it requires the training data to conform to the distribution of $\pi_{\text{ref}}$.
Figure 9: Comparison of using eq. 11 and eq. 13 as target.
When $\pi_{\text{ref}}$ fails to capture previous successful experiences, it becomes difficult to sample these experiences from $\pi_{\text{ref}}$, which in turn, can lead to further forgetting of those experiences. (2) Both eq. 11 and eq. 13 aim to approximate the distance between two distributions. The use of Mean Squared Error as a metric to measure the distance between policy distributions is also employed in Fujimoto & Gu (2021). (3) Eq. 13 is unable to decrease the probability of actions with negative advantage. When correct action is hard to sample from $\pi_{\text{ref}}$, eq. 13 may instead increase the probability of actions with negative advantage, compounding the challenge of sampling correct actions. We also conduct experiments to further validate the superiority of eq. 11, with the results presented in Figure 9. In these experiments, eq. 13 is trained exclusively on data sampled from $\pi_{\text{ref}}$, while eq. 11 utilize both replay buffer data and sampled data.
When discount factor $\gamma$ is used. When considering the discount factor, eq. 9 is modified as follows:
$\beta\log\frac{\pi^{*}(a_{t}|s_{t},I)}{\pi_{\text{ref}}(a_{t}|s_{t},I)}=r(s_{t},a_{t},I)+\gamma V^{*}(s_{t+1},I)-V^{*}(s_{t})=A^{*}(s_{t},a_{t},I)$ (14)
This will not affect the subsequent derivation of the policy loss function (eq. 12). The loss function for the critic remains unchanged. However, the equation to calculate the advantage is modified as follows:
$A(s_{t},\!a_{t},\!I)\!=\!\lambda\big{(}r(s_{t},\!a_{t},\!I)+\gamma V(s_{t+1},\!I)\!-\!V(s_{t},\!I)\big{)}\!+\!(1\!-\!\lambda)\big{(}\gamma^{T-t}r(s_{T},\!a_{T},\!I)\!-\!V(s_{t},\!I)\big{)}$ (15)
In our experiment, we use the discount factor and set its value to 0.9.
We use eq. 7 and eq. 8 to estimate $A^{*}(s,a,I)$. The training of critic (eq. 7) also operates in an off-policy manner.
A.2 Proof
To prove the feasibility of replacing $A^{*}$ with an advantage estimator $A$ based on a value network $V$, which is trained on data collected throughout the training process, we aim to demonstrate that using $V$ to train the policy can lead to policy improvement.
The current policy is denoted by $\pi_{\text{old}}$. The training dataset $D$ is composed of trajectories sampled using $\pi_{\text{old}}$ as well as historical experiences stored in a replay buffer. The behavioral policy associated with the dataset $D$ is referred to as $\pi_{\mu}$. The corresponding value function and action-value function under $\pi_{\mu}$ are denoted as $V^{\pi_{\mu}}$ and $Q^{\pi_{\mu}}$, respectively. Below, we demonstrate that minimizing the objective $\arg\min_{\theta}\mathbb{E}_{D}[(\log_{\pi_{\theta}}(a_{t}|s_{t})-\exp(\frac{Q^{\pi_{\mu}}(s_{t},a_{t})-V^{\pi_{\mu}}(s_{t})}{\beta}))^{2}]$ can lead to policy improvement. For brevity, we omit $I$ in the following proof.
Let the new policy $\pi_{\text{new}}(a_{t}|s_{t})$ be defined as:
$\pi_{\text{new}}(a_{t}|s_{t})=\exp\left(\frac{Q^{\pi_{\mu}}(s_{t},a_{t})-V^{\pi_{\mu}}(s_{t})}{\beta}\right)$ (16)
It can be shown that: $D_{\text{KL}}(\pi_{\text{new}}||\pi_{\text{new}})\leq D_{\text{KL}}(\pi_{\mu}||\pi_{\text{new}})$. From this, we derive the following inequality:
$\begin{split}&\mathbb{E}_{a_{t}\sim\pi_{\text{new}}}\!\left[\log\pi_{\text{new}}(a_{t}|s_{t})\!-\!\frac{Q^{\pi_{\mu}}(s_{t},a_{t})}{\beta}\!+\!\frac{V^{\pi_{\mu}}(s_{t})}{\beta}\right]\\ \leq&\ \mathbb{E}_{a_{t}\sim\pi_{\mu}}\!\left[\log\pi_{\mu}(a_{t}|s_{t})\!-\!\frac{Q^{\pi_{\mu}}(s_{t},a_{t})}{\beta}\!+\!\frac{V^{\pi_{\mu}}(s_{t})}{\beta}\right]\end{split}$ (17)
Since $V^{\pi_{\mu}}(s_{t})$ depends only on the state $s_{t}$ and not on the actions, it can be factored out of the expectation. We can rewrite the inequality as:
$\mathbb{E}_{a_{t}\sim\pi_{\text{new}}}\left[Q^{\pi_{\mu}}(s_{t},a_{t})-\beta\log\pi_{\text{new}}(a_{t}|s_{t})\right]\geq V^{\pi_{\mu}}(s_{t})$ (18)
where $V^{\pi_{\mu}}(s_{t})=\mathbb{E}_{a\sim\pi_{\mu}(a_{t}|s_{t})}[Q^{\pi_{\mu}}(s_{t},a_{t})-\beta\log\pi_{\mu}(a_{t}|s_{t})]$, which is satisfied in maximum entropy reinforcement learning setting (Haarnoja et al., 2018).
Consider the soft Bellman equation:
$\begin{split}Q^{\pi_{\mu}}(s_{t},a_{t})&=r(s_{t}|a_{t})+\beta\log\pi_{\text{ref}}(a_{t}|s_{t})+V^{\pi_{\mu}}(s_{t+1})\\ &\leq r(s_{t}|a_{t})+\beta\log\pi_{\text{ref}}(a_{t}|s_{t})+\mathbb{E}_{a_{t+1}\sim\pi_{\text{new}}}[Q^{\pi_{\mu}}(s_{t+1},a_{t+1})-\beta\log\pi_{\text{new}}(a_{t+1}|s_{t+1})]\\ &\cdots\\ &\leq Q^{\pi_{\text{new}}}(s_{t},a_{t})\end{split}$ (19)
where we iteratively expand $Q^{\pi_{\mu}}$ by applying the soft Bellman equation and the bound provided in Eq 18. Since $\pi_{\text{new}}$ is an improved policy over $\pi_{\mu}$ (i.e., $Q^{\pi_{\text{new}}(s_{t},a_{t})}\geq Q^{\pi_{\mu}}(s_{t},a_{t})$), and $\pi_{\theta}$ is trained to closely match $\pi_{\text{new}}$, it follows that $\pi_{\theta}$ will also be an improved policy over $\pi_{\mu}$ provided the approximation is accurate. Formally, we have:
$V^{\pi_{\theta}}(s)\approx V^{\pi_{\text{new}}(s)}\geq V^{\pi_{\mu}}(s)$ (20)
Since we only incorporate the correct trajectory from the historical data, we can assert that policy $\pi_{\mu}$ will not perform worse than the old policy $\pi_{\text{old}}$. Therefore, we have the relationship:
$V^{\pi_{\theta}}(s)\approx V^{\pi_{\text{new}}}(s)\geq V^{\pi_{\mu}}(s)\geq V^{\pi_{\text{old}}}(s)$ (21)
The policy improvement process is well-established. By iteratively performing policy evaluation and policy improvement, the policy $\pi_{\theta}$ can be made to converge toward the optimal policy $\pi^{*}$. If the data in the replay buffer is not utilized, the same theoretical process holds. However, in this case, $V^{\pi_{\mu}}$ and $Q^{\pi_{\mu}}$ are replaced by $V^{\pi_{\text{old}}}$ and $Q^{\pi_{\text{old}}}$.
Figure 10: Comparison of using $V^{\pi_{\mu}}$ and $V^{\pi_{\text{old}}}$.
We conduct an experiment to evaluate the effectiveness of using $V^{\pi_{\mu}}$ instead of $V^{\pi_{\text{old}}}$ in WebRL training. In this experiment, the policy $\pi_{\theta}$ is trained using data from two sources: rollout trajectories during the current phase and past experiences stored in the replay buffer. For the critic, we compare two training settings: training with data from both rollout trajectories and experiences in the replay buffer versus using only rollout trajectories. The results, as presented in Figure 10, reveal that the model performs better when the critic is trained with data from both the replay buffer and the rollout trajectories. This observation underscores the significance of integrating experiences from prior successful trajectories into the critic’s training process, which can lead to more effective policy improvement.
A.3 Comparison
The comparison result between the policy update algorithm of WebRL and other algorithms is shown in Table 4.
Table 4: Comparison between WebRL and different algorithms based on various criteria.
Algorithm Off-policy Pair-wise Data KL-constrained Target Reduce Error Prob
WebRL [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DPO [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PPO [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AWR [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Comparison with DPO (Rafailov et al., 2024b). DPO also solves for the optimal value of eq. 1, obtaining the relationship between the optimal policy $\pi^{*}$ and the reward $r$. Rather than explicitly solving for the reward $r$, DPO employs contrastive learning by constructing paired data to build the learning target of $\pi_{\theta}$. This approach allows DPO to bypass the need to estimate $r$. In contrast, our approach explicitly fits the value function $V$ to guide the optimization of $\pi_{\theta}$, making it converge towards the optimal policy $\pi^{*}$. However, DPO requires pair-wise data, which is challenging to obtain in web tasks. This is primarily because implementing state backtracking on web pages is difficult, making it hard to collect outcomes of different actions on the same page state.
Comparison with PPO (Schulman et al., 2017). Eq.1 is also the target of RLHF. Prior work often employs PPO to optimize this objective. However, as an on-policy algorithm, PPO has low sampling efficiency and requires large amounts of data for stable improvement. This limitation makes it unsuitable for environments such as WebArena and real-world websites, where interaction is expensive and inefficient. Instead, our approach operates in an off-policy manner.
Comparison with AWR (Peng et al., 2019). AWR is one of the widely used offline RL algorithms. The loss function of AWR is $\operatorname*{arg\,max}_{\theta}\mathbb{E}_{\nu}\left[\log\pi_{\theta}(a|s,I)\cdot\exp(A(s,a,I)/\beta\right]$. Compared to our method, AWR has a different optimization goal. Our approach explicitly minimizes the KL divergence between the reference policy $\pi_{\text{ref}}$ and the current policy $\pi_{\theta}$, while AWR focuses on constraining the KL divergence between the current policy and the behavioral policy derived from training data $\nu$ (instead of the reference policy $\pi_{\text{ref}}$). Additionally, AWR does not directly reduce the probability of incorrect actions. In web tasks, sampled data often include unsuccessful traces. Ignoring such data wastes valuable information, while using it risks increasing the likelihood of incorrect actions. In contrast, our method effectively reduces the probability of incorrect actions.
Appendix B Training Details
B.1 Details of WebRL
The agent observation consists of three components:
  • $\bullet$
    User instruction: The instruction provided by the user.
  • $\bullet$
    Action history: A record of the actions the agent has previously taken.
  • $\bullet$
    Webpage HTML: HTML content of the current webpage.
We process the HTML, simplifying its structure and assigning distinct element IDs to all clickable elements. This facilitates the model’s ability to identify and indicate which specific element requires manipulation.
The agent actions are mostly similar to those defined in WebArena-Lite (Liu et al., 2024b), with some additional actions.
  • $\bullet$
    Click: Clicks an element with a specific ID.
  • $\bullet$
    Hover: Hovers over an element with a specific ID.
  • $\bullet$
    Type: Types a message into an input box with a specific ID.
  • $\bullet$
    Search: Types a message into an input box with a specific ID and presses Enter to initiate a search.
  • $\bullet$
    Press: Emulates a specific keyboard key combination.
  • $\bullet$
    Scroll: Scrolls the page up or down.
  • $\bullet$
    Select_dropdown_option: Selects an option from a dropdown menu with a specific ID.
  • $\bullet$
    New_tab: Opens a new tab in the current browser.
  • $\bullet$
    Tab_focus: Switches focus to a browser tab at a specified index.
  • $\bullet$
    Close_tab: Closes the current tab.
  • $\bullet$
    Goto: Navigates to a specific URL.
  • $\bullet$
    Go_back: Returns to the previous page.
  • $\bullet$
    Go_forward: Moves to the next page if available.
  • $\bullet$
    Exit: Terminates the operation, returns the response, and exits.